Country Switzerland (Franc CHF Euro € BID (EUR/CHF)=0.65/1.25; Ask (EUR/CHF)=0.68/1.15 USD equivalent BID ASK $0.65/CHF $0.68please show all steps
Answer #1

Given that, $/CHF = Bid 0.65 and Ask 0.68

and, $/€ = Bid 1.15 and Ask 1.2

Therefore, €/CHF will be: Bid = Bid €/$ × Bid $/CHF = (1/Ask $/€) × Bid $/CHF = (1/1.2)×0.65 = 0.5417

(Note: There seems to be a typo mistake in the question. In the table, Ask $/€ is 1.2, whereas in Bid(EUR/CHF) it is 0.65/1.25. Here, the solution is given, considering the table rate i.e. 1.2 as correct.)

And €/CHF Ask will be: Ask €/$ × Ask $/CHF = (1/Bid $/€) × Ask $/CHF = (1/1.15)×0.68 = 0.5913

From above, i.e. from €/CHF, CHF/€ will be calculated as follows:

Bid CHF/€ = 1/(Ask €/CHF) = 1/0.5913 = 1.6912

Ask CHF/€ = 1/(Bid €/CHF) = 1/0.5417 = 1.846

(If this was helful, kindly rate the answer positively. Thank You:)